OpenStudy (anonymous):

Please give an example of a function with more than one horizontal asymptote.

OpenStudy (lgbasallote):

do you want a graph? or an equation?

OpenStudy (anonymous):

An equation please!

OpenStudy (lgbasallote):

ugh. i can only do graphs. sorry

OpenStudy (anonymous):

its okay :)

OpenStudy (anonymous):

\[\frac{ 2 }{ 1+e ^{x} }-1\]

OpenStudy (anonymous):

Does that work?

OpenStudy (unklerhaukus):

\[y^2=\frac{x+1}{x}\]

OpenStudy (anonymous):

okay could you explain why those work?

OpenStudy (anonymous):